# Modeling, Inference, and Prediction in Mobility-Based Compartmental Models for Epidemiology

### v1

@@ 3.1 Theoretical guarantees
Let $I(m,t)$ be the infected mobility distribution of the mobility-based SIRS model (4) with the initial condition (6), yielding the 0th-order moment $Q_{0}^{I}(t)$. Assume that $Q_{0}^{I}(m,t)$ is arbitrarily smooth (i.e., $Q_{0}^{I}\in\mathcal{C}^{\infty}(0,\infty)$) and all the parameters ($I_{0}$, $\beta$, $\gamma$, and $\varepsilon$) are fixed and not equal to 0, then there is a one-to-one correspondence between $f$ and $Q_{0}^{I}$.
This theorem indicates that for any two different mobility distributions ($f\neq\tilde{f}$), the corresponding time series of the infected numbers ($Q_{0}^{I}$ and $Q_{0}^{\tilde{I}}$) also differ. This implies that it is sufficient to infer $f$ from $I$ theoretically and this inference problem has a unique solution. We prove this theorem at the end of this subsection.
Recall the moment notation in (18). We multiply $m^{k}$ to (4) and obtain the equations for moments
$\frac{\mathrm{d}}{\mathrm{d}t}Q_{k}^{S}$ $=-\beta Q_{1}^{I}Q_{k+1}^{S}+\varepsilon Q_{k}^{R}$ (42a)
$\frac{\mathrm{d}}{\mathrm{d}t}Q_{k}^{I}$ $=\beta Q_{1}^{I}Q_{k+1}^{S}-\gamma Q_{k}^{I}$ (42b)
@@
Consider another set of solutions $\tilde{S}(m,t),\tilde{I}(m,t),\tilde{R}(m,t)$ of (4) with the initial condition proportional to $\tilde{f}(m)$.
@@
$Q_{k}^{S}=Q_{k}^{\tilde{S}},\quad Q_{k}^{I}=Q_{k}^{\tilde{I}},\quad\text{and}\quad Q_{k}^{R}=Q_{k}^{\tilde{R}}.$ (43)
@@
which utilizes the fact that $Q_{k}^{f}=Q_{k}^{S}+Q_{k}^{I}+Q_{k}^{R}$. Similarly, we have
@@
Step 1. We first show that $Q_{k}^{I}=Q_{k}^{\tilde{I}}$ is equivalent to $Q_{k}^{S}=Q_{k}^{\tilde{S}}$. Suppose that $Q_{k}^{I}=Q_{k}^{\tilde{I}}$. We compute the difference between (44) and (45) and obtain
@@
Since $Q_{k}^{I}(0)=I_{0}Q_{k}^{f}$ and $Q_{k}^{\tilde{I}}(0)=I_{0}Q_{k}^{\tilde{f}}$, we use $Q_{k}^{I}(0)=Q_{k}^{\tilde{I}}(0)$ and obtain
$Q_{k}^{f}=Q_{k}^{\tilde{f}},\quad Q_{k}^{S}(0)-Q_{k}^{\tilde{S}}(0)=S_{0}\left(Q_{k}^{f}-Q_{k}^{\tilde{f}}\right)=0.$ (47)
@@
Step 2. We then show that $Q_{k}^{I}=Q_{k}^{\tilde{I}}$ is equivalent to $Q_{k}^{R}=Q_{k}^{\tilde{R}}$. Suppose that $Q_{k}^{R}=Q_{k}^{\tilde{R}}$. Therefore, we have $\mathrm{d}Q_{k}^{R}/\mathrm{d}t=\mathrm{d}Q_{k}^{\tilde{R}}/\mathrm{d}t$. From (42c), we have $Q_{k}^{I}=Q_{k}^{\tilde{I}}$. Suppose $Q_{k}^{I}=Q_{k}^{\tilde{I}}$. Using the similar idea as above, we have $e(t)=Q_{k}^{R}-Q_{k}^{\tilde{R}}$ satisfies
@@
Theorem 5 and Lemma 6 indicate that there is a one-to-one correspondence between $f$ and $Q_{0}^{S}$ and that between $f$ and $Q_{0}^{R}$. This implies that, alternatively, we can also infer the mobility distribution $f$ from the time series of $Q_{0}^{S}$ or $Q_{0}^{R}$ if they are available and the inference problems yield unique solutions.
@@
Let $L$ be a positive integer and the moments satisfy
@@
$\frac{\mathrm{d}^{\ell}}{\mathrm{d}t^{\ell}}Q_{k}^{S,I,R}(0)$ $\neq\frac{\mathrm{d}^{\ell}}{\mathrm{d}t^{\ell}}Q_{k}^{\tilde{S},\tilde{I},\tilde{R}}(0),\quad k=K-\ell.$
@@
We use the same moment notation (18). Since $f\neq\tilde{f}$, we can find a positive integer $K$ such that
@@
We plug the initial values into (42) and obtain
@@
Notice that (58) satisfies the conditions in Lemma 7 for $L=2$. By induction, we know (51) hold for $\ell=2,3,\ldots,K$. In particular, we use the inequality result in Lemma 7 for $\ell=K$ and $k=0$ yielding
@@
3.2 Learning the mobility distribution using a neural-network framework
Mobility distribution significantly impacts the pandemic size and transmission. However, in reality, it is not feasible to directly measure the mobility density functions of different compartments during pandemics. Realistically, the available data in epidemics is often limited to the number of infected individuals. Theorem 5 suggests a one-to-one correspondence between the mobility distribution $f(m)$ and the ratio of the infected population $\langle I\rangle(t)$. Thus, a natural approach is to infer the mobility distribution from the early-stage pandemic data to make accurate predictions for the future course of the pandemic. We propose a neural-network-based approach to learn the operator from the time-series $\langle I\rangle(\cdot)$ to the mobility distribution $f(\cdot)$.
Assume that the pandemic data takes the form of $\mathbf{I}=(\langle I\rangle(t_{1}),\ldots,\langle I\rangle(t_{N_{t}}))^{T}$, which is a time series of the ratio of the infected population at discrete time steps. We discretize $\mathbf{f}$ using the discrete cosine transform (DCF) and denote the first $N_{f}$ modes as $\mathbf{f}=(\hat{f}_{1},\ldots,\hat{f}_{N_{f}})^{T}$. We consider a feedforward neural network $\mathrm{G}_{\theta}$ to learn the mapping between $\mathbf{I}$ and $\mathbf{f}$:
@@ 3.2 Learning the mobility distribution using a neural-network framework
We sample the mobility distribution $f$ and obtain $\langle I\rangle$ by simulating the mobility-based SIRS model (4) to form the training set. When using neural networks to learn unknown parameters or functions, it is crucial to sample a training set that encompasses a wide range of scenarios. In our context, this involves properly sampling the mobility distribution comprehensively and robustly. Since a Gaussian mixture model is a universal approximator of probability densities, we sample $f$ by a Gaussian mixture with random coefficients. In particular, we let
${f}\propto\sum_{n=1}^{3}r_{n}\mathcal{N}(\mu_{n},\sigma_{n}^{2}),$ (62)
where $\mathcal{N}(\mu,\sigma^{2})$ is the Gaussian density function with mean $\mu$ and variance $\sigma^{2}$ and $r_{n}$, $\mu_{n}$, and $\sigma_{n}$ are all uniformly distributed on $[0,1]$, $[-0.2,1.2]$, and $[0.01,0.26]$, respectively.
Suppose the dataset $\mathcal{D}$ contains $K$ pairs of $\mathbf{I}$ and $\mathbf{f}$, which we obtain by discretizing $\langle I\rangle$ and $f$. We parametrize $\mathrm{G}_{\theta}$ with a fully-connected feedforward neural network and minimize the loss function
$\mathrm{L}(\theta)=\frac{1}{K}\sum_{i=1}^{K}\left[\mathrm{G}_{\theta}(\mathbf{I}_{i})-\mathbf{f}_{i}\right]^{2}.$
In our simulations, we use a six-layer-fully connected neural network. The width of layers are 128, 512, 512, 512, 128, and 32, respectively. Each layer includes a Sigmoid activation function and $L^{2}$ regularization with a magnitude of $0.005$. In addition, we choose $N_{f}=8$ and $N_{t}=101$. For each training step, we sample a random batch from the training set and use the Adam optimizer with a learning rate of 0.0005 to minimize the loss function for each batch. We generate a training set of size 10,000 and train the neural network for 2,000 epochs. During each training epoch, we optimize the loss function using 1,562 batches with a batch size of 54.
3.3 Numerical results
@@
Given the dataset $\mathcal{D}$, we obtain the mapping $\mathrm{G}_{\theta}$ by minimizing the loss function (3.2). To evaluate performance, we generate an additional six mobility distributions using (62) and simulate the corresponding time series of infected populations $\mathbf{I}$ as inputs. Figure 6 shows a comparison between the learned mobility distributions and the ground truth. It is noteworthy that the testing inputs $\mathbf{I}_{i}$ are not part of the original training set; however, the inferred mobility distributions are capable of capturing the trends of the ground truth and yield time series that closely match the ground truth.
(a) Predicted and true mobility distributions $f(m)$
(b) Predicted and true infected populations $\langle I\rangle(t)$
Figure 6: (a): A comparison between the predicted mobility distributions and the true distributions. (b): A comparison between the predicted and true $\langle I\rangle(t)$.
@@
We analyze mobility distributions using real COVID-19 data, shown in Figure 7. From February 29th, 2020, to September 16th, 2020, covering the initial 200 days of our dataset, we examine the infection ratio between Massachusetts and New York. We calculate the total infected cases by summing daily new infections up to the current time, adjusted with a decaying prefactor to account for previously infected individuals. Specifically, we use
$\langle I\rangle(t)=I_{\text{daily}}(t)+\exp{(-1/\gamma)}\langle I\rangle(t-1).$ (63)
We recover the daily true case counts $I_{\text{daily}}(t)$ from the death data and fit them with exponential growth for the first 30 days. The real data reflects a much smaller scale of disease transmission compared to the training set. To use the real data as input, we scale the input by a factor of 50. We show the resulting input data in Figure 7(a) and display the learned mobility distributions in Figure 7(b).
@@
Figure 7: (a): The scaled ratio of COVID-19 infected cases during the initial wave in Massachusetts and New York from February 29th, 2020, to September 16th, 2020. (b): Inferred mobility distributions corresponding to the inputs in (a).
@@ 4 Conclusion and outlook
In this paper, we integrate social heterogeneity into compartmental models by introducing a mobility variable $m$, assuming it ranges from 00 to $1$ for each individual. This approach represents the population in each compartment (susceptible, infected, recovered, etc.) as a density function over $[0,1]$. Individuals with higher mobility are more prone to transmitting infectious diseases to others. The time evolution of population density for each compartment is described through an integro-differential equation.
We explore the key properties of the proposed mobility-based compartmental models crucial for epidemiological modeling, such as the basic reproduction number $\mathcal{R}_{0}$, final pandemic size, and mean mobility. The proposed mobility-based compartment models also effectively mitigate the overestimation problem inherent in many homogeneous compartmental models when estimating the final pandemic size. The basic reproduction rate $\mathcal{R}_{0}$ is often estimated based on the initial growth rate of the infected population. In the early stages of a pandemic, individuals with higher mobility tend to get infected quickly, contributing to a higher $\mathcal{R}_{0}$. As the pandemic evolves, the recovered population with higher mobility can impart some degree of immunity to the broader population, particularly those with lower mobility. This process helps in mitigating the eventual final size of the pandemic. We rigorously demonstrate that under the same $\mathcal{R}_{0}$, the classical SIR model yields the largest final pandemic size. Numerical findings consistently indicate that a polarized mobility distribution typically results in a significantly smaller final pandemic size compared to the classical SIR model. Additionally, we establish a one-to-one correspondence between the time series of the total infected population and the mobility distribution, demonstrating that the mobility distribution can be practically inferred from infection data. We propose a deep learning-based algorithm for inferring the mobility distribution directly from the total infected population.
The primary objective of integrating mobility into compartmental disease models is to utilize the mobility distribution as an intermediary connecting disease dynamics with social dynamics. Changes in public health policies and shifts in public opinion about infectious diseases influence disease spread by altering people’s mobility patterns. Moving forward, we plan to incorporate opinion dynamics into mobility-based compartmental models to explore the interplay between public opinion and disease transmission dynamics.

### v2

@@ 3.1 Identifiability of the mobility distribution
Let $I(m,t)$ be the infected mobility distribution of the mobility-based SIRS model (4) with the initial condition (6) and $Q_{0}^{I}(t)$ be the 0th order moment of $I(m,t)$, defined in (42). Assume that $Q_{0}^{I}(t)$ is arbitrarily smooth (i.e., $Q_{0}^{I}(t)\in\mathcal{C}^{\infty}(0,\infty)$) and that all the parameters ($I_{0}$, $\beta$, $\gamma$, and $\varepsilon$) are fixed and not equal to 0, then there is a one-to-one correspondence between $f$ and $Q_{0}^{I}$.
Recall the moment notation in (17). We multiply $m^{k}$ to (4) and integrate over $m$, obtaining the equations for moments
$\frac{\mathrm{d}}{\mathrm{d}t}Q_{k}^{S}$ $=-\beta Q_{1}^{I}Q_{k+1}^{S}+\varepsilon Q_{k}^{R},$ (43a)
$\frac{\mathrm{d}}{\mathrm{d}t}Q_{k}^{I}$ $=\beta Q_{1}^{I}Q_{k+1}^{S}-\gamma Q_{k}^{I},$ (43b)
@@
Let $\tilde{S}(m,t),\tilde{I}(m,t),\tilde{R}(m,t)$ be the mobility distributions of (4) with the initial condition
$\tilde{S}(m,0)=(1-I_{0})\tilde{f}(m),\quad\tilde{I}(m,0)=I_{0}\tilde{f}(m),\quad\tilde{R}(m,0)=0,$ (44)
and $Q_{k}^{\tilde{S}},Q_{k}^{\tilde{I}},Q_{k}^{\tilde{R}}$ be their $k$-th order moments.
@@
$Q_{k}^{S}(t)=Q_{k}^{\tilde{S}}(t),\quad$ for all $\text{for all\leavevmode\nobreak\ }t\geq 0,$ (45)
$Q_{k}^{I}(t)=Q_{k}^{\tilde{I}}(t),\quad$ for all $\text{for all\leavevmode\nobreak\ }t\geq 0,$
$Q_{k}^{R}(t)=Q_{k}^{\tilde{R}}(t),\quad$ for all $\text{for all\leavevmode\nobreak\ }t\geq 0.$
@@
which uses the fact that $Q_{k}^{f}=Q_{k}^{S}+Q_{k}^{I}+Q_{k}^{R}$. Similarly, we have
@@
Step 1. We show that $Q_{k}^{I}=Q_{k}^{\tilde{I}}$ is equivalent to $Q_{k}^{S}=Q_{k}^{\tilde{S}}$. Suppose that $Q_{k}^{I}=Q_{k}^{\tilde{I}}$. We compute the difference between (46) and (47) and obtain
@@
Since $Q_{k}^{I}(0)=I_{0}Q_{k}^{f}$ and $Q_{k}^{\tilde{I}}(0)=I_{0}Q_{k}^{\tilde{f}}$, we use $Q_{k}^{I}(0)=Q_{k}^{\tilde{I}}(0)$ and obtain $Q_{k}^{f}=Q_{k}^{\tilde{f}}$, so that
$Q_{k}^{S}(0)-Q_{k}^{\tilde{S}}(0)=(1-I_{0})\left(Q_{k}^{f}-Q_{k}^{\tilde{f}}\right)=0.$ (49)
@@
Step 2. We then show that $Q_{k}^{I}=Q_{k}^{\tilde{I}}$ is equivalent to $Q_{k}^{R}=Q_{k}^{\tilde{R}}$. Suppose that $Q_{k}^{R}=Q_{k}^{\tilde{R}}$, then $\mathrm{d}Q_{k}^{R}/\mathrm{d}t=\mathrm{d}Q_{k}^{\tilde{R}}/\mathrm{d}t$. From (43c), we have $Q_{k}^{I}=Q_{k}^{\tilde{I}}$. Suppose $Q_{k}^{I}=Q_{k}^{\tilde{I}}$. Using the similar idea as above, we have $e(t)=Q_{k}^{R}-Q_{k}^{\tilde{R}}$ satisfies
@@
Theorem 6 and Lemma 7 indicate that $f(m)$ also has a one-to-one correspondence to $Q_{0}^{S}(t)$ and $Q_{0}^{R}(t)$. This implies that, alternatively, we can infer the mobility distribution $f(m)$ from $Q_{0}^{S}(t)$ or $Q_{0}^{R}(t)$ if they are available, and the inference problems have unique solutions.
@@
Let $L$ and $K$ be positive integers and $K\geq 2$. Assume the moments satisfy
@@
$\frac{\mathrm{d}^{\ell}}{\mathrm{d}t^{\ell}}Q_{k}^{S,I,R}(0)$ $\neq\frac{\mathrm{d}^{\ell}}{\mathrm{d}t^{\ell}}Q_{k}^{\tilde{S},\tilde{I},\tilde{R}}(0),\quad k=K-\ell,$
@@
We use the same moment notation (17). Assume that $f\neq\tilde{f}$, then we can find a positive integer $K$ such that
@@
We plug the initial values of moments into (43) and obtain
@@
Notice that (60) satisfies the conditions in Lemma 8 for $L=2$. By induction, we know (53) hold for $\ell=2,3,\ldots,K$. In particular, we use the inequality result in Lemma 8 for $\ell=K$ and $k=0$ and obtain
@@
3.2 Numerical inference of the mobility distribution
Theorem 6 establishes a one-to-one correspondence between the mobility distribution $f(m)$ and the infected population ratio $Q_{0}^{I}(t)$. This provides a theoretical basis for inferring the mobility distribution from time series data of the infected population. One can infer the mobility distribution from early-stage pandemic data and use it to predict the future course of the pandemic. In this section, we propose a numerical approach using neural networks to achieve this inference.
3.2.1 A neural-network-based learning framework
Let $\langle I\rangle(t)$ be the ratio of the infected population at time $t$ (i.e., $\langle I\rangle(t)=\int_{0}^{1}I(m,t)\leavevmode\nobreak\ \mathrm{d}m$) and $\mathbf{I}$ be the observed time series of $\langle I\rangle(t)$, which takes the form of
$\mathbf{I}=(\langle I\rangle(t_{1}),\langle I\rangle(t_{2}),\ldots,\langle I\rangle(t_{n_{t}}))^{T}.$ (63)
Since the data is a finite-dimensional vector, we are only able to infer the mobility distribution $f$ in a finite-dimensional space. Let $\bar{f}$ be the spatial discretization of $f$ with $M$ uniform grids on the interval $[0,1]$. We take the discrete cosine transform (DCT) of $\bar{f}$ and denote the DCT of $\bar{f}$ by $\hat{f}=\sqrt{M}(1,\hat{f}_{1},\ldots,\hat{f}_{n_{f}},\ldots)$. Let $\mathbf{f}$ be the first $n_{f}$ nontrivial modes of $\hat{f}$, i.e.,
$\mathbf{f}=(\hat{f}_{1},\hat{f}_{2},\ldots,\hat{f}_{n_{f}})^{T}.$ (64)
We formulate the inference problem using a feedforward neural network $\mathrm{G}_{\theta}$ to learn the mapping from $\mathbf{I}$ to $\mathbf{f}$:
@@ 3.2.1 A neural-network-based learning framework
To build the training set, we sample various mobility distributions $f(m)$ and simulate the mobility-based SIRS model (4) to generate the corresponding time series of $\langle I\rangle(t)$. When using neural networks to learn unknown parameters or functions, it’s essential to sample a training set that covers a wide range of scenarios, ensuring the model is effective across the prediction regime. In our context, this involves sampling the mobility distribution comprehensively and robustly. Since the Gaussian mixture model is a universal approximator of probability densities [13], we sample $f$ as a Gaussian mixture model with random parameters. In particular, we let
${f}\propto\sum_{i=1}^{\alpha}w_{i}\mathcal{N}(\mu_{i},\sigma_{i}^{2})$ (66)
where $\mathcal{N}(\mu,\sigma^{2})$ is the Gaussian probability density function and $\mu_{i}$, $\sigma_{i}$, and $w_{i}$ are random numbers. Suppose the dataset contains $K$ pairs of $\mathbf{I}_{k}$ and $\mathbf{f}_{k}$, which we discretize from $\langle I_{k}\rangle(t)$ and $f_{k}(m)$. Then, we determine the fully connected feedforward neural network $\mathrm{G}_{\theta}$ by minimizing the loss function $\mathrm{L}(\theta)$:
$\mathrm{L}(\theta)=\frac{1}{K}\sum_{k=1}^{K}\|\mathrm{G}_{\theta}(\mathbf{I}_{k})-\mathbf{f}_{k}\|^{2},$ (67)
@@
3.2.2 Numerical results
In our simulations, we let $f_{k}$ be a Gaussian mixture of 3 normal distributions:
${f_{k}}\propto\sum_{i=1}^{3}w_{i}\mathcal{N}(\mu_{i},\sigma_{i}^{2}),$ (68)
and sample the parameters $w_{i}$, $\mu_{i}$, and $\sigma_{i}$ as random numbers uniformly on intervals $[0,1]$, $[-0.1,1.1]$, and $[0.05,1.05]$, respectively. We solve the mobility-based SIRS model (4) for $\langle I\rangle(t)$ with the initial condition (6) and parameters $\beta=0.8,\varepsilon=0.005$, $\gamma=0.14$, and $I_{0}=0.0001$. We sample $K=10,000$ mobility distributions $f_{k}$ in the form of (68) and compute corresponding $\langle I\rangle(t)$. We discretize $f_{k}$ and $\langle I\rangle(t)$ with $n_{t}=101$ in (63) and $n_{f}=8$ in (64) and obtain the training set containing $K=10,000$ input-output pairs $(\mathbf{I}_{k},\mathbf{f}_{k})$. To improve the performance, we scale the input data $\mathbf{I}_{k}$ by a factor of $50$ and the output data $\mathbf{f}_{k}$ by a factor of $3$. The mean values of the input and output sets are around $1$ after rescaling.
In our simulations, we use a six-layer fully connected neural network, and the width of layers are 128, 512, 512, 512, 128, and 32, respectively. Each layer uses a Sigmoid activation function and an $L^{2}$ regularization of $\theta$ with a magnitude of $0.005$. For each training step, we sample a random batch from the training set and use the Adam optimizer with a learning rate of 0.0005 to minimize the loss function for each batch. We train the neural network for 2,000 epochs, and during each training epoch, we optimize the loss function using 1,562 batches with a batch size of 54.
@@ 3.2.2 Numerical results
We obtain the mapping $\mathrm{G}_{\theta}$ by minimizing the loss function (67). To test the performance of the learned mapping $\mathrm{G}_{\theta}$, we generate 6 additional mobility distributions ($\mathbf{f}_{1},\ldots,\mathbf{f}_{6}$) using (68) and simulate the corresponding time series of infected populations ($\mathbf{I}_{1},\ldots,\mathbf{I}_{6}$) as inputs. We compare the predicted outputs $\hat{\mathbf{f}}_{i}=\mathrm{G}_{\theta}(\mathbf{I}_{i})$ for $i=1,\ldots,6$ with the ground truth $\mathbf{f}_{i}$. We compute their inverse DCTs and show the comparison between the learned mobility distributions $\hat{f}_{i}$ and the actual distributions ${f}_{i}$ in Figure 6. The learned distributions successfully capture the overall trends of the true mobility distributions.
Figure 6: A comparison of 6 predicted mobility densities $\hat{f}_{i}(m)$ and their ground truth $f_{i}(m)$.
We solve the mobility-based SIRS model (4) using the predicted mobility $\hat{{f}}_{i}$ and obtain the corresponding time series of the infected population $\langle\hat{I}_{i}\rangle(t)$. Figure 7 compares these predicted time series with the ground truth. The results show that the inferred mobility distributions generate time series of the infected population that closely match the ground truth.
Figure 7: A comparison of the predicted infected time series $\hat{\langle I_{i}\rangle}(t)$ and their ground truth $\langle I_{i}\rangle(t)$.
@@
Next, we focus on inferring mobility distributions using real-world COVID-19 data [15]. The original dataset includes daily new infection counts for Massachusetts and New York from February 29th to September 16th, 2020. We calculate the total number of infected cases by summing the daily new infections with the previously infected cases, adjusted by a decaying factor to account for the recovered cases. In particular, we assume that the infected population $\langle I\rangle(t)$ follows
$\langle I\rangle(t)=I_{\text{daily}}(t)+\exp{(-\gamma)}\langle I\rangle(t-1),$ (69)
where $I_{daily}(t)$ is the daily infection cases and $\gamma=0.1345$ is the recovery rate, obtained from fitting an SIR model with the Massachusetts data obtained in [15]. We scale the input $\langle I\rangle(t)$ by a factor of $50$, which is similar to the treatment of the training set. We show the rescaled time series of the COVID-19 infected population in Figure 8, together with the inferred mobility distributions from this dataset.
The real-world data is influenced by numerous factors that have drastically altered social mobility during this period, such as social distancing and lockdown policies. Although our model does not account for these factors, the inferred mobility distribution still provides valuable insights. We observe that the inferred distributions include a small portion of high-mobility individuals, which accelerates disease transmission during the early stages of the COVID-19 pandemic. Classical models, which assume a homogeneous population, often fail to distinguish these high-mobility groups, leading to an overestimation of the pandemic size.
@@
Figure 8: (a) The COVID-19 infection ratio (scaled by $50$) during the initial wave in Massachusetts and New York from February 29th, 2020, to September 16th, 2020. (b) Inferred mobility distributions from the data in (a).
@@ 4 Conclusion and outlook
We introduce and analyze density-based compartmental models in epidemiology that incorporate population heterogeneity through mobility. We assume that each individual has a mobility variable ranging from 0 to 1 and that individuals with higher mobility are more likely to transmit infectious diseases. We describe the disease dynamics by modeling the evolution of mobility distribution functions of the three compartments: susceptible, infected, and recovered. Notably, the classical SIR and SIRS models are special cases of our model when the mobility distributions are Dirac delta functions.
We derive key epidemiological parameters for our mobility-based compartmental models, including the basic reproduction number $\mathcal{R}_{0}$, the final pandemic size, and mean mobility. Our models effectively address the overestimation issue common in many homogeneous compartmental models when estimating the final pandemic size. Early in a pandemic, individuals with higher mobility tend to become infected more rapidly, resulting in a higher initial growth rate of cases compared to a homogeneous population. As the pandemic evolves, the recovered high-mobility population imparts immunity to the broader population, particularly those with lower mobility. Most existing models neglect this mechanism, while ours captures it, offering a more accurate estimate of the basic reproduction number and the final pandemic size. We prove that the classical SIR model, where the entire population has the same mobility (i.e., mobility distributions are Dirac delta functions), yields the largest final pandemic size compared to models with other mobility distributions. Numerical results show that a polarized mobility distribution results in a significantly smaller final pandemic size than the classical SIR model.
As it is often challenging to measure mobility distributions in practice, we formulate an inference problem to estimate the mobility distribution from the ratio of the total infected population. We prove that there is a one-to-one correspondence between the time series of the total infected population and the mobility distribution, which implies that the mobility distribution can be uniquely identified from the infection data. Additionally, we propose a numerical approach using a machine learning framework to infer the mobility distribution from the data and validate our results with real COVID-19 data.
In this paper, we incorporate heterogeneous mobility distributions into the SIR and SIRS models. It is also valuable to integrate mobility aspects into more complex compartmental models, such as those with additional compartments or those accounting for birth and death. Currently, we assume that each individual’s mobility is time-independent, but one can relax this assumption to consider mobility dynamics influenced by social opinions, public policies, overall infection cases, and many other factors. Another potentially interesting extension is to combine the mobility-based compartmental models with control problems and explore how managing population mobility can reduce disease spread.